OpenStudy (sloppycanada):

Write the equation of a parabola with a vertex at (-5, 2) and a directrix y = -1.

OpenStudy (sloppycanada):

\[x+5 = \frac{ 1 }{ 12 }(y-2)^2\]

OpenStudy (amoodarya):

directix is y=-1 so parabola is vertical (x+5)^2=4p(y-2)

OpenStudy (sloppycanada):

Oh so I find 'p' now?

OpenStudy (anonymous):

Do you have any question choices? sorry haven't done these in a long time lol

OpenStudy (sloppycanada):

http://gyazo.com/92c75a607eefcd67e67bf08e7368c6fa

OpenStudy (anonymous):

I remember now! So we first put it in vertex form. Typical vertex form is ;\[y=a(x-h)^2+k\]

OpenStudy (anonymous):

Keep in mind that h is x, and k is y, so \[(h,k) = (x,y)\]

OpenStudy (anonymous):

this seems hard lol

OpenStudy (anonymous):

So we have the vertex (and forgot to mention, that h,k is the vertex) (-5,2)

OpenStudy (sloppycanada):

Yup! And since it's not at (0,0) it means I have to add things. If I remember correct, it's the opposite sign that you're supposed to use.

OpenStudy (anonymous):

we put them in, and now we get what?

OpenStudy (sloppycanada):

x+5 = (y-2)^2

OpenStudy (sloppycanada):

But I still need the 4p

OpenStudy (anonymous):

\[y=a(x+5)^2-2\]

OpenStudy (anonymous):

Well we still have to find a, and note the answer choice that you picked is wrong :/

OpenStudy (sloppycanada):

Which is something to do with "p"

OpenStudy (anonymous):

hold on a sec...

OpenStudy (sloppycanada):

Okay...?

OpenStudy (anonymous):

SO a is the slope

OpenStudy (anonymous):

Or the the rate of change

OpenStudy (loser66):

@amoodaray gave you the equation, why do you go on the complex way?

OpenStudy (sloppycanada):

I'm not sure how to find 'p'

OpenStudy (loser66):

Is it not that "directrix y =-1" given?? and it gives you p = 1?? am I right?

OpenStudy (loser66):